In performing the following dihybrid testcross, I obtain the following F2 phenotypes. What is the distance between the loci that control flower color and pollen length? P: Long, red x white, short, F1: Long, red x white, short, F2: 450 Red, long, 470 white, short, 42 Red, short, 38 white, long
a. 0.8 cM
b. 80 cM
c. 8 cM
d. 10 cM
e. 1 cM
Answer #1
  • Recombination frequency (RF) indicates the frequency of chromosomal crossover between two genes.
  • RF measures genetic linkage and the loci of gene on a chromosome.
  • Recombination frequency (measured in %) determines the distance of two gene loci (measured in centi-Morgan (cM) or as map unit (MU).
  • During recombination process (like meiotic recombination, where recombination occurs between non-sister chromatids), the further the genes are the more is their tendency to cross over or higher is the recombination frequency (as indicated by Morgan).

RF= Number of recombinants/ (Number of recombinants + Number of non-recombinants) x 100

In given case:

Recombinants = 42 + 38 = 80

Total = 450+470+42+38 =1000

RF = 80 /1000 x 100 = 8 = map distance = 8 cM